[Dsg-hallc_controls] CAEN-HV boards test results

Marc Mcmullen mcmullen at jlab.org
Tue Jul 30 11:36:13 EDT 2019

Hi Simona and Brad,

DSG is investigating the issue Pablo described in the previous email, where certain channel parameters of a CAEN HV module will arbitrarily change. The tests Pablo ran were on the A1535 modules, which were tested in slot 13. All of these modules were old, and it was not clear if the issue was local to that particular type of module or the slot.

I am currently testing the brand new A7030TN modules on multiple slots. Module 0297 is located in slot 1, channel 28 failed to complete a full current test by ramping to ~1500V with a 1.5M ohm load, it stopped at 400V. I checked the parameters using the CAEN GECO interface and the SVMax was set to 400V.

Prior to running the test I had set all slot 1 SVMax values to 3000V.

Before rerunning the test this morning I set the SVMax back to 3000V (using the GECO interface) and ran the test. The test stopped at 600V, I checked the SVMax and it was set to 3000V. At that point I asked Pablo to check the PV using CSS to look directly into the server and the SVMax process variable was indeed set to 600V, the GECO interface was still displaying 3000V.

I used the GECO interface and re-entered 3000V and noted that as I highlighted the value, the GECO display changed to 400V. I changed the value to 3000V and re-ran the test and it ramped to the expected voltage before and tripped as expected.

This confirms a suspicion that the embedded EPICS server software in the sy4527 mainframe has an issue. The process variable changed multiple times to different values. We need to test this on another sy4527 crate to see if this is a systemic problem or just a bad crate. When it is convenient please send us the other crate and we will continue 7030TN testing, and see if the issue manifests itself on the second crate.
Note in the pics below that channel 01.028 SVMax is set to 3000 V in the top screen, which is the GECO interface, while below CSS is displaying 600 V for the same variable.



From: Dsg-hallc_controls <dsg-hallc_controls-bounces at jlab.org> on behalf of Pablo Campero Rojas <campero at jlab.org>
Sent: Thursday, July 25, 2019 1:49 PM
To: dsg-hallc controls <dsg-hallc_controls at jlab.org>
Cc: Patrizia Rossi <rossi at jlab.org>
Subject: [Dsg-hallc_controls] CAEN-HV boards test results

Hello Simona,

We developed an EPICS based program and CSS screens to test the three CAEN-HV boards that brought to the DSG-lab last week.

From the test performed, we found that all three boards have CAEN software issues and one of them has a hardware issue as well.

  Software issues:

  1.  Some channels’ parameters like “Imon”, “Vset”, VRdwn”, “Imax”, “Vmon” arbitrarily change from its pre-set values (set values at test time) to random values.
  2.  Discrepancies between GECO-2020 (CAEN Control software) and the PVs values (Generated by CAEN EPICS Server) displayed for some of the above mentioned variables.

 However, we were able to ramp all channels to the set voltage after we solved these CAEN software issues by re-entering set values, re-setting power On/Off  channels, and sometimes resetting the built-in CAEN EPICS Server.

  Hardware issue:

  1.  Channel 22 of HV board (S/N: 556) does not to ramp up.

We recommend sending HV board S/N: 556 back to CAEN.

Attached below is a table with the summary of the findings for each board tested. To see detailed results from the test, you are most than welcome to come and visit us in the DSG room EEL-121.

Best Regards,


Board #

Board Model

Serial Number

Hardware Issue








For some channels, Vset stuck at a different value than the pre-set value during the test.  One of the channels tripped, because "Imax" value changed from its set value during test.

Discrepancies between GECO-2020 values and EPICS PVs values for Imon.

All affected channels ramped to desired set value after the set values were re-entered and power On/Off reset.

For the recovery of some channels, reset of the built-in CAEN EPICS Server was required.





For some channels, Vset value changed from its pre-set value.

For variable Vmon, found discrepancies between GECO-2020 values and EPICS PVs values.

All affected channels ramped to desired set value after the set values were re-entered and channels were reset.






Set values for Vset, VRdwn and Vmax changed (PVs and value in GECO-2020) from its pre-set values during the test.

Re-entering values for Vset, VRdwn and Vmax, then reset power On/Off allowed ramp up all channels to the desired set voltage except for channel 22, which did not ramp up.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/dsg-hallc_controls/attachments/20190730/e6fe7c16/attachment-0001.html>
-------------- next part --------------
A non-text attachment was scrubbed...
Name: ch28 svmax.JPG
Type: image/jpeg
Size: 194985 bytes
Desc: ch28 svmax.JPG
URL: <https://mailman.jlab.org/pipermail/dsg-hallc_controls/attachments/20190730/e6fe7c16/attachment-0001.jpe>
-------------- next part --------------
A non-text attachment was scrubbed...
Name: HVCAEN-A7030TN-297-ch28-SVMax Issue.png
Type: image/png
Size: 399331 bytes
Desc: HVCAEN-A7030TN-297-ch28-SVMax Issue.png
URL: <https://mailman.jlab.org/pipermail/dsg-hallc_controls/attachments/20190730/e6fe7c16/attachment-0001.png>

More information about the Dsg-hallc_controls mailing list